# Image Intrinsic Scale Assessment: Bridging the Gap Between Quality and Resolution

### v1

@@ Abstract
Image Quality Assessment (IQA) measures and predicts perceived image quality by human observers. Although recent studies have highlighted the critical influence that variations in the scale of an image have on its perceived quality, this relationship has not been systematically quantified. To bridge this gap, we introduce the Image Intrinsic Scale (IIS), defined as the largest scale where an image exhibits its highest perceived quality. We also present the Image Intrinsic Scale Assessment (IISA) task, which involves subjectively measuring and predicting the IIS based on human judgments. We develop a subjective annotation methodology and create the IISA-DB dataset, comprising 785 image-IIS pairs annotated by experts in a rigorously controlled crowdsourcing study. Furthermore, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a strategy that leverages how the IIS of an image varies with downscaling to generate weak labels. Experiments show that applying WIISA during the training of several IQA methods adapted for IISA consistently improves the performance compared to using only ground-truth labels. We will release the code, dataset, and pre-trained models upon acceptance.
@@
Figure 1: Top: downscaling an image affects its perceived quality. As the original image (left) is downscaled, degradation becomes less noticeable, but some high-frequency details may be lost. The optimal trade-off occurs at the image’s intrinsic scale. Note that the first two images are cropped for clearer visualization. Bottom: we plot several possible profiles, corresponding to different images, depicting how quality changes with scale. The Image Intrinsic Scale (IIS) is the largest scale at which an image shows its highest perceived quality. The IIS is measured above a lower bound $s_{lb}\!=\!0.05$, where the image quality is more interpretable.
@@ 1 Introduction
Most IQA datasets annotate quality ratings – in the form of Mean Opinion Score (MOS) – for images at a fixed resolution [14, 15, 9]. To the best of our knowledge, KonX [34] was the first dataset that provided quality annotations for each image rescaled across multiple resolutions. The authors showed that perceived quality varies with image resolution (i.e., size in pixels). When an image is downscaled, its resolution decreases. Since images are typically displayed at a consistent pixel density on any given display medium, downscaled images also appear physically smaller. The perceived quality of smaller images often increases because degradation becomes less noticeable, e.g., noise grain becomes smaller or fades. Conversely, relevant high-frequency details may get lost when the resolution is too low, leading to relatively lower perceived quality. The top section of Fig. 1 shows an example of the former phenomenon. The image at the original size ($scale=1$) appears blurry and degraded. However, when downscaled ($scale=0.25$), the degradations are not noticeable, improving quality, while relevant details, such as the bird’s feathers and barbs, are crisp and clear. At smaller scales ($scale=0.1$), the individual barbs cannot be resolved anymore, and detail is lost.
Based on the earlier observation, we aim to answer the question: what is an image’s “true” resolution? Or, in other words, what is the optimal scale of an image that provides the best balance between visibility of degradation and retention of relevant high-frequency details? Thus, we introduce the concept of Image Intrinsic Scale (IIS), which is the largest scale at which an image shows its highest perceived quality. We select the largest scale as it maximizes the availability of high-quality information. Moreover, we propose a new task named Image Intrinsic Scale Assessment (IISA). IISA aims to subjectively measure and predict the IIS in alignment with human perception. Our main goal is to develop effective automated methods for predicting the IIS, and this first requires establishing a reliable annotation methodology to collect subjective opinions.
IISA has multiple potential applications in both industry and research: 1) it optimizes image display and storage by identifying an image’s peak quality across scales, crucial for applications like web, print, and gaming. This enhances user experience and saves storage or printing space by using the display media’s pixel density (e.g., PPI) to determine optimal display sizes. 2) IISA improves quality measurement sensitivity over traditional NR-IQA methods through implicit cross-scale comparative judgments, making it effective for detecting subtle quality degradations from restoration methods like denoising and super-resolution. See the supplementary material for more details on sensitivity. 3) IISA supports the creation of better low-level vision datasets. Unlike current methods that reduce diversity by discarding low-quality images, by downscaling high-resolution images to their IIS, IISA can help maintain both quality and diversity, essential for diverse and high-quality training datasets [12, 23].
@@
To automatically predict the IIS, we initially assess the zero-shot transfer capabilities of pre-trained NR-IQA models. The subpar performance of these models on IISA underscores the distinct nature of the two tasks and the need for IISA-specific training.
To this end, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a straightforward yet effective approach for generating multiple weakly labeled image-IIS pairs from a single ground-truth one. WIISA leverages a unique property of the intrinsic scale that allows us the extrapolate the ground-truth IIS of an image to its downscaled versions via a piece-wise function. Our method can be easily integrated into the training of any model, as it is independent of specific training strategies or model characteristics. The experiments demonstrate that the proposed approach consistently boosts the performance of several methods from the NR-IQA domain retrained for IISA on the proposed dataset.
@@
  • We design an annotation methodology for collecting subjective opinions of the IIS that ensures high participant self-consistency and reliability of the labels. Moreover, we create the IISA-DB dataset, comprising 785 image-IIS pairs annotated by image quality experts in a thoroughly controlled crowdsourcing study.
@@ 2 Related Work
Another line of research has focused on the effects of upscaling on image quality.
Works such as [29, 38] aim to discriminate between authentic and artificially upscaled images. While valuable for detecting upscaling artifacts, such a binary classification does not offer a generalizable image quality metric.
Kansy et al. [19] introduce the concept of ‘effective resolution”, i.e., the smallest size an image can be downscaled to, such that when it is upscaled back to its original dimensions the loss of detail is not perceptible. This is distinct from the IIS. Consider the example of a noisy image: the scale of its effective resolution should be high enough to preserve the noise pattern such that upscaling will not lead to noticeable detail loss relative to the original. In contrast, the IIS of the image should be low enough to make the noise not visible anymore, improving the perceived quality of the image.
@@ 3 Image Intrinsic Scale Assessment
As illustrated in Fig. 1 and supported by previous studies [34, 13, 8, 25, 11], the relationship between the perceived quality of an image and its presentation scale – defined as the scale at which it is displayed for subjective annotation – is generally assumed to follow a concave-down or monotonic function for scales smaller than 1 (i.e., the original size). When an image is downscaled, quality initially improves as degradation like blur and noise decrease. However, downsizing past a certain point reduces perceived quality primarily due to loss of detail and interpolation artifacts like aliasing.
@@
In our implementation we set a lower bound $s_{lb}=0.05$ for the scale $s$, because reliably evaluating the quality of very small images is challenging due to the inherent loss of detail from downscaling. When assessing image quality, slight variations in size typically result in minimal quality change. Therefore, an image might be perceived as having its highest quality at multiple scales. Thus, we take the maximum value among these scales in Eq. 1, as it maximizes the presence of high-quality information.
@@ 4.1 Dataset collection
We collected 20 opinions per image from 10 freelance participants, meeting the minimum of 15 per stimulus recommended by [17] for quality assessment. Given the satisfactory inter-group agreement shown in Fig. 4, the number of participants is adequate. This aligns with similar IQA studies like SPAQ [9] and UHD-IQA [15], which gather 15 and 20 opinions per image, respectively. All freelancers had previous experience annotating IQA datasets and a professional background in visual arts, such as photography and graphic design. Before starting the main task, participants had to undergo a training phase. Specifically, they annotated images with predefined ground truth IIS ranges determined by the authors of this paper, with textual hints provided to guide accurate annotation. After the training stage, participants annotated images presented in randomly ordered batches of 90. Each batch was viewed twice, with a few days between repetitions. We measured Spearman’s Rank-order Correlation Coefficient (SRCC) between the annotations of repetitions of corresponding batches for each participant. Pairs of batches with an SRCC lower than 0.5 were required to be re-annotated. This process ensures accurate individual annotations and an overall high experiment reliability.
@@
Based on our annotation strategy, we created the IISA-DB dataset, comprising 785 image-IIS pairs labeled at their original resolution via crowdsourcing. All images are larger than 2048 pixels in width (see Fig. 3). We started from an initial set of 900 authentically distorted real-world images. We sampled 300 of them from the KonIQ-10k [14] IQA dataset according to their MOS and machine tags to ensure a diverse range of quality levels and content types. The remaining 600 images were sourced from the Pixabay website and selected for diversity based on multiple factors, such as user tags, number of likes, and downloads. Finally, for privacy reasons, we manually reviewed the dataset to remove elements that could identify people, resulting in the final set of 785 images. Due to the sampling procedure, our dataset contains images with diverse subjects, quality levels, resolutions, and consequently, IIS values.
Figure 3: Distribution of the width of the images of the IISA-DB dataset at their original and intrinsic scale.
@@
We obtained the ground truth IIS labels for each image (i.e. $\Omega(I)$) by aggregating individual annotations from all participants, similar to how Mean Opinion Scores (MOS) are calculated for image quality ratings. Therefore, we refer to these aggregated labels as the Mean Opinion Intrinsic Scale (MOIS). This aggregation effectively reduces the variability in perceived quality due to the diverse viewing conditions inherent in crowdsourcing. Considering the non-linear relationship between scale values in our slider UI – where equal ratios between two scale values are represented by different lengths on the slider – we chose the geometric mean to compute the MOIS. For more details on the impact of viewing conditions on annotation reliability, refer to the supplementary material.
@@
4.2 Dataset Analysis
Across the 785 images composing our dataset, the average MOIS is 0.347, with per-image MOIS ranging from 0.060 to 0.811. Fig. 3 illustrates the distribution of image widths at their original and intrinsic scales. The noticeable gap between these distributions highlights how quality generally increases by downscaling.
To assess the reliability of the subjective annotations of IISA-DB, we measure the inter-group agreement between different participant sub-groups and compare it with that of existing NR-IQA datasets [15, 34, 14]. Specifically, for datasets where each participant annotated each image twice (IISA-DB, KonX [34], UHD-IQA [15]), we randomly sample without replacement pairs of non-overlapping groups of up to five participants each, thus considering at most ten annotations per image. For KonIQ-10k [14], since participants provided just a single opinion per image and each rated only a subset of images, we sample the required number of opinions per group without considering participant identity. For each dataset, we sample 200 pairs of groups and compute the average SRCC and RMSD (Root Mean Squared Difference) between the MOIS (IISA) or MOS (NR-IQA) of each group. Note that although both IIS and quality ratings are scaled to the $[0,1]$ range, their absolute errors are not directly comparable due to the different nature of their scales – non-linear and linear, respectively. Therefore, SRCC is a more appropriate measure than RMSD to compare inter-group agreement. Nonetheless, we report both for completeness in Fig. 4. We observe that the proposed IISA-DB dataset has SRCC values that are comparable to existing reliable NR-IQA datasets, such as KonIQ-10k. The same applies to the RMSD. Overall, these results demonstrate the reliability of the annotations in our dataset.
@@ 5.1 Pre-trained NR-IQA Models
Method SRCC PLCC RMSE MAE
TOPIQ (SPAQ) 0.475 0.437 0.136 0.104
TOPIQ (UHD-IQA) 0.587 0.600 0.382 0.364
@@ 6.2 Results
6.3 Discussion
Proposed method
Tab. 3 shows the results. Regarding the number of weak labels, the experiments show that increasing the value of $n_{wl}$ leads to better results up to $n_{wl}=2$, but beyond that performance declines. We hypothesize that too many weak labels introduce overly redundant information, as the image content remains constant across them, with only the scale varying. Likewise, a high downscaling threshold leads to the weak labels being overly similar to the ground-truth ones, thus reducing the performance improvement. Conversely, lower $\delta$ values allow the weakly labeled IIS to be higher (see Eq. 2), possibly resulting in them being out of the ground-truth distribution (the average IIS in IISA-DB is 0.347). Finally, we observe that using the bilinear and bicubic interpolation algorithms for generating the weak labels leads to marginally lower performance than using the Lanczos one. This finding suggests that the interpolation algorithm has limited impact in training models with WIISA.
Assumptions
Figure 6: Average quality of the images belonging to the KonX dataset across three resolutions. The images are divided into three groups, based on the resolution for which the quality is maximum.
In Sec. 3 we assume that the relationship between image quality and scale follows a concave-down or monotonic function. To verify this, we consider the KonX dataset [34], which provides quality annotations for images at three different resolutions. Similarly to [8], we divide the images into three groups based on the resolution for which their MOS is maximum and compute the average quality scores. We plot the results in Fig. 6. We observe that the quality of all three groups varies according to either a concave-down or monotonic function, confirming that our assumption holds on average. On an individual level, we find that 90% of the images (378 out of 420) show MOS trends across resolutions that are consistent with our assumption. We attribute the instances where the assumption does not hold to potential noise in the annotation process, due to the presence of interpolation artifacts or subjective biases from annotators. See the supplementary material for more details. Furthermore, studies that vary viewing distances while keeping image resolution fixed [25, 11, 8] also support our assumption. Indeed, they show that image quality follows a concave-down function with respect to viewing distance. Together, these results provide strong evidence that our assumption holds in most cases.

### v2

@@ Abstract
Image Quality Assessment (IQA) measures and predicts perceived image quality by human observers. Although recent studies have highlighted the critical influence that variations in the scale of an image have on its perceived quality, this relationship has not been systematically quantified. To bridge this gap, we introduce the Image Intrinsic Scale (IIS), defined as the largest scale where an image exhibits its highest perceived quality. We also present the Image Intrinsic Scale Assessment (IISA) task, which involves subjectively measuring and predicting the IIS based on human judgments. We develop a subjective annotation methodology and create the IISA-DB dataset, comprising 785 image-IIS pairs annotated by experts in a rigorously controlled crowdsourcing study with verified reliability. Furthermore, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a strategy that leverages how the IIS of an image varies with downscaling to generate weak labels. Experiments show that applying WIISA during the training of several IQA methods adapted for IISA consistently improves the performance compared to using only ground-truth labels. We will release the code, dataset, and pre-trained models upon acceptance.
@@
Figure 1: Top: downscaling an image affects its perceived quality. As the original image (left) is downscaled, degradation becomes less noticeable, but some high-frequency details may be lost. The optimal trade-off occurs at the image’s intrinsic scale. Note that the first two images are cropped for clearer visualization. Bottom: we plot several possible profiles corresponding to different images, depicting how the quality changes with scale – quality refers to the latent expected value of measured subjective quality ratings rather than the noisy mean observed from finite samples. The Image Intrinsic Scale (IIS) is the largest scale at which an image shows its highest perceived quality. The IIS is measured above a lower bound $s_{lb}\!=\!0.05$, where the image quality is more interpretable.
@@ 1 Introduction
Most IQA datasets annotate quality ratings – in the form of Mean Opinion Scores (MOS) – for images at a fixed resolution [14, 15, 9]. To the best of our knowledge, KonX [33] was the first dataset that provided quality annotations for each image rescaled across multiple resolutions. The authors showed that perceived quality varies with image resolution (i.e., size in pixels). When an image is downscaled, its resolution decreases. Since images are typically displayed at a consistent pixel density on any given display medium, downscaled images also appear physically smaller. The perceived quality of smaller images often increases because degradation becomes less noticeable, e.g., noise grain becomes smaller or fades. Conversely, relevant high-frequency details may get lost when the resolution is too low, leading to relatively lower perceived quality. The top section of Fig. 1 shows an example of the former phenomenon. The image at the original size ($scale=1$) appears blurry and degraded. However, when downscaled ($scale=0.25$), the degradations are not noticeable, improving quality, while relevant details, such as the bird’s feathers and barbs, are crisp and clear. At smaller scales ($scale=0.1$), the individual barbs cannot be resolved anymore, and detail is lost.
Based on the earlier observation, we aim to answer the question: what is the optimal scale of an image that provides the best balance between visibility of degradation and retention of relevant high-frequency details? Thus, we introduce the concept of Image Intrinsic Scale (IIS), which is the largest scale at which an image shows its highest perceived quality. We select the largest scale as it maximizes the availability of high-quality information. Moreover, we propose a new task named Image Intrinsic Scale Assessment (IISA). IISA aims to subjectively measure and predict the IIS in alignment with human perception. Our main goal is to develop effective automated methods for predicting the IIS, and this first requires establishing a reliable annotation methodology to collect subjective opinions.
Subjective studies measure latent variables that are not directly observable – here, the perceived quality or IIS. These latent variables are commonly modeled as random variables, with each human opinion representing a sample. The theoretical “true” value is the expectation of this random variable. In principle, when we refer to image quality or IIS, we discuss the expected value, which can be estimated by averaging many independent and unbiased opinions, e.g., via the MOS. In practice, the number of opinions is limited by cost and other constraints, so we work with approximations that depend on careful experimental design to minimize bias and achieve reliable estimates.
@@
To automatically predict the IIS, we initially assess the zero-shot transfer capabilities of pre-trained NR-IQA models. The subpar performance of these models on IISA underscores the distinct nature of the two tasks and the need for IISA-specific training. To this end, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a straightforward yet effective approach for generating multiple weakly labeled image-IIS pairs from a single ground-truth one. WIISA leverages a unique property of the intrinsic scale that allows us the extrapolate the ground-truth IIS of an image to its downscaled versions via a piece-wise function. Our method can be easily integrated into the training of any model, as it is independent of specific training strategies or model characteristics. The experiments demonstrate that the proposed approach consistently boosts the performance of several methods from the NR-IQA domain retrained for IISA on the proposed dataset.
@@
  • We design an annotation methodology for collecting subjective opinions of the IIS that ensures high participant self-consistency and reliability of the labels. Moreover, we create the IISA-DB dataset, comprising 785 image-IIS pairs annotated by image quality experts in a thoroughly controlled and reliable crowdsourcing study.
@@ 2 Related Work
Another line of research has focused on the effects of upscaling on image quality. Works such as [28, 37] aim to discriminate between authentic and artificially upscaled images. While valuable for detecting upscaling artifacts, such a binary classification does not offer a generalizable image quality metric. Kansy et al. [18] introduce the concept of ‘effective resolution”, i.e., the smallest size an image can be downscaled to, such that when it is upscaled back to its original dimensions the loss of detail is not perceptible. This is distinct from the IIS. Consider the example of a noisy image: the scale of its effective resolution should be high enough to preserve the noise pattern such that upscaling will not lead to noticeable detail loss relative to the original. In contrast, the IIS of the image should be low enough to make the noise not visible anymore, improving the perceived quality of the image.
@@ 3 Image Intrinsic Scale Assessment
As illustrated in the bottom section of Fig. 1 and supported by previous studies [33, 13, 8, 24, 11], the relationship between the perceived quality of an image and its presentation scale – defined as the scale at which it is displayed for subjective annotation – is generally assumed to follow a concave-down or monotonic function for scales smaller than 1 (i.e., the original size). When an image is downscaled, quality initially improves as degradation like blur and noise decrease. However, downsizing past a certain point reduces perceived quality primarily due to loss of detail and interpolation artifacts like aliasing.
@@
In our implementation, we set a lower bound $s_{lb}=0.05$ for the scale $s$ because reliably evaluating the quality of very small images is challenging due to the inherent loss of detail from downscaling. When assessing image quality, slight variations in size typically result in minimal quality change. Therefore, an image might be perceived as having its highest quality at multiple scales. Thus, we take the maximum value among these scales in Eq. 1, as it maximizes the presence of high-quality information.
@@ 2. A2:
Applications:
IISA addresses applications where understanding scale-dependent quality is critical: 1) it optimizes image display and storage by identifying an image’s peak quality across scales, crucial for applications like the web, print, and gaming. This enhances user experience and saves storage or printing space by using the display media’s pixel density (e.g., PPI) to determine optimal display sizes. 2) IISA improves quality measurement sensitivity over traditional NR-IQA methods through implicit cross-scale comparative judgments, making it effective for detecting subtle quality degradations from restoration methods like denoising and super-resolution. See the supplementary material for more details on sensitivity. 3) IISA supports the creation of better low-level vision datasets. Unlike current methods that reduce diversity by discarding low-quality images, by downscaling high-resolution images to their IIS, IISA can help maintain both quality and diversity, essential for diverse and high-quality training datasets [12, 22]. See the supplementary for more details on the applications of IISA.
@@ 4.1 Dataset collection
We collected 20 opinions per image from 10 freelance participants, meeting the minimum of 15 per stimulus recommended by [16] for quality assessment. Given the satisfactory confidence intervals reported in Sec. 4.2, the number of participants is adequate. This aligns with similar IQA studies like SPAQ [9] and UHD-IQA [15], which gather 15 and 20 opinions per image, respectively. All freelancers had previous experience annotating IQA datasets and a professional background in visual arts, such as photography and graphic design. Before starting the main task, participants had to undergo a training phase. Specifically, they annotated images with predefined ground truth IIS ranges determined by the authors of this paper, with textual hints provided to guide accurate annotation. After the training stage, participants annotated images presented in randomly ordered batches of 90. Each batch was viewed twice, with a few days between repetitions. We measured Spearman’s Rank-order Correlation Coefficient (SRCC) between the annotations of repetitions of corresponding batches for each participant. Pairs of batches with an SRCC lower than 0.5 were required to be re-annotated. This process ensures accurate individual annotations and an overall high experiment reliability.
@@
Following our annotation strategy, we created the IISA-DB dataset, comprising 785 image-IIS pairs labeled at their original resolution via crowdsourcing. All images are larger than 2048 pixels in width. We started from an initial set of 900 authentically distorted real-world images. We sampled 300 of them from the KonIQ-10k [14] IQA dataset according to their MOS and machine tags to ensure a diverse range of quality levels and content types. The remaining 600 images were sourced from the Pixabay website and selected for diversity based on multiple factors, such as user tags and number of likes. Finally, to address privacy concerns, we manually removed any images containing identifiable individuals, resulting in the final set of 785 images. As a result of the sampling procedure, our dataset contains images with diverse subjects, quality levels, resolutions, and consequently, IIS values.
@@
We obtained the ground truth IIS labels for each image (i.e. $\Omega(I)$) by aggregating individual annotations from all participants, similar to how Mean Opinion Scores (MOS) are calculated for image quality ratings. Therefore, we refer to these aggregated labels as the Mean Opinion Intrinsic Scale (MOIS). This aggregation effectively reduces the variability in perceived quality due to the diverse viewing conditions inherent in crowdsourcing. Considering the non-linear relationship between scale values in our slider UI – where different lengths on the slider represent equal ratios between two scale values – we chose the geometric mean to compute the MOIS.
4.2 Dataset Analysis
Reliability of the annotations
To assess the reliability of the subjective annotations of IISA-DB, we compute the confidence intervals (CIs) for the MOIS of each image. Following a well-established methodology, we sample 20 opinions with replacement for each image, compute their geometric mean, and repeat the process 100 times. The 95% CI is then estimated as half the range between the $2.5^{\text{th}}$ and $97.5^{\text{th}}$ percentiles of the simulated geometric means. This yields an average CI of $0.057$ for our dataset. As a reference, we consider the KonX NR-IQA dataset, whose annotation conditions were similar to ours, involving remote studies with freelance participants. The average CI of the MOS for the KonX dataset is $0.046$. Given the differences between the IISA and NR-IQA tasks, some variation in the CIs is expected. Therefore, this outcome indicates that our annotation process has a level of reliability comparable to a highly reliable NR-IQA dataset such as KonX.
@@ 4.2 Dataset Analysis
We further assess the reliability of our annotations using a common approach that measures the inter-group agreement between different participant sub-groups and compares it with that of existing NR-IQA datasets [15, 33, 14]. Specifically, for datasets where each participant annotated each image twice (IISA-DB, KonX [33], UHD-IQA [15]), we randomly sample without replacement pairs of non-overlapping groups of up to five participants each, thus considering at most ten annotations per image. For KonIQ-10k [14], since participants provided just a single opinion per image and each rated only a subset of images, we sample the required number of opinions per group without considering participant identity. For each dataset, we sample 200 pairs of groups and compute the average SRCC and RMSD (Root Mean Squared Difference) between the MOIS (IISA) or MOS (NR-IQA) of each group. Note that although both IIS and quality ratings are scaled to the $[0,1]$ range, their absolute errors are not directly comparable due to the different nature of their scales – non-linear and linear, respectively. Therefore, SRCC is more appropriate than RMSD in comparing inter-group agreements. Nonetheless, we report both for completeness in Fig. 3. We observe that the proposed IISA-DB dataset has SRCC values that are comparable to existing reliable NR-IQA datasets, such as KonIQ-10k. The same applies to the RMSD. Overall, these results demonstrate the reliability of our dataset’s annotations.
Effects of variable viewing conditions
Viewing conditions – including display characteristics (contrast, brightness), environmental lighting, viewing distance, and viewer acuity – significantly influence perceived image quality. Laboratory experiments tightly control these factors but consequently measure atypical conditions with limited variability, restricting the applicability of the findings. In contrast, crowdsourcing involves diverse and less controlled conditions, potentially reducing individual participant reliability. Nevertheless, when reliable group opinions are obtained through well-controlled methods, the resulting assessments generalize better. Our annotation framework has demonstrated high reliability, ensuring that despite individual variation, group aggregates accurately measure the perceptual IIS. Thus, we provide highly reliable data that better reflects real-world viewing conditions, enhancing the validity and applicability of our findings.
Method SRCC PLCC RMSE MAE
TOPIQ (SPAQ) 0.475 0.437 0.136 0.104
TOPIQ (UHD-IQA) 0.587 0.600 0.382 0.364
@@ 6.2 Results
6.3 Discussion
Proposed method
Tab. 3 shows the results. Regarding the number of weak labels, the experiments show that increasing the value of $n_{wl}$ leads to better results up to $n_{wl}=2$, but beyond that performance declines. We hypothesize that too many weak labels introduce overly redundant information, as the image content remains constant across them, with only the scale varying. Likewise, a high downscaling threshold leads to the weak labels being overly similar to the ground-truth ones, thus reducing the performance improvement. Conversely, lower $\delta$ values allow the weakly labeled IIS to be higher (see Eq. 2), possibly resulting in them being out of the ground-truth distribution (the average IIS in IISA-DB is 0.347). Finally, we observe that using the bilinear and bicubic interpolation algorithms to generate the weak labels leads to marginally lower performance than using the Lanczos one. This finding suggests that the interpolation algorithm has limited impact in training models with WIISA.
@@ 6.3 Discussion
Assumption
In Sec. 3 we assume that the relationship between image quality and scale follows a concave-down or monotonic function. To verify this, we consider the KonX dataset [33], which provides quality annotations for images downscaled to three predefined resolutions. We find that 90% of the image triplets (378 out of 420) show MOS trends across resolutions that are consistent with our assumption. We attribute the instances where the assumption does not hold to potential noise in the annotation process caused by subjective biases from annotators or interpolation artifacts. See the supplementary material for more details. Furthermore, studies that vary viewing distances while keeping image resolution fixed [24, 11, 8] also support our assumption. They show that image quality follows a concave-down function with respect to viewing distance. In principle, changing the viewing distance at a fixed image resolution is equivalent to changing the resolution (and display size) at a fixed viewing distance. Together, these studies provide compelling evidence supporting the validity of our assumption across a wide range of conditions.